OpenStudy (anonymous):

write an equation of a line in point slope form that has a slope of -3 and passes through the point 3,-4.

OpenStudy (anonymous):

Point slope form looks like: \[y-y _{1}=m(x-x _{1})\] where m=slope and \[(x _{1},y _{1})\] is a point on the line. All you need it to plug in what you were given. \[y-(-4)=-3(x-3)=y+4=3(x-3) \] Hope this helps!

OpenStudy (anonymous):

so y+4=3(x−3) is the equation?o:

OpenStudy (anonymous):

Yes that's the equation in point slope form. If you wanted it in slope intercept form you would basically subtract 4 from both sides and then simplify but since you want it in point slope form then yes. TYPO: I just realized I accidentally put in 3 instead of -3.... Should really be: \[y+4=-3(x-3)\] You can see your point (3, -4) and the slope of -3

OpenStudy (anonymous):

so then if i were to write another equation for a line in point slope form that has a slope of -3/4 and passes through the poin 4,5. the equation would be y-5=-3/4(x−4)?

OpenStudy (anonymous):

Yep! You got it! I had to look up which form was which because you will mostly use slope-intercept form in the future but yes you are right!!!

OpenStudy (anonymous):

slope intercept form is y=mx+b right?

OpenStudy (anonymous):

Yep! Point-slope form gives you more information in the equation but slope-intercept form gives you the only two pieces of information you really need in my opinion.

OpenStudy (anonymous):

alrighty thank you so much :D

OpenStudy (anonymous):

No problem!